# RelCon: Relative Contrastive Learning for a Motion Foundation Model for Wearable Data

### v4

@@ 5.3 Ablation Study Results
The Importance of Accel Augmentations: Invariances can capture conceptually important information for motion data, and augmentations have been used extensively in prior approaches towards capturing these invariances (Yurtman & Barshan, 2017; Florentino-Liaño et al., 2012; Zhong & Deng, 2014). For example, learning 3D rotation invariance allows the embeddings to be robust to how different users may wear or carry an accelerometry device. In our ablation study in Table 5, the w/o Augmentations results were generated by removing the augmentations from the training of the learnable distance measure. Removing augmentations had a strong negative impact on results, across tasks. RelCon uses augmentations to learn a motif-similarity based distance metric that is used with a relative loss. While prior approaches using augmentations learn only invariances (or variances, in the case of augmentation prediction), RelCon incorporates learned invariances along with motifs and within-person dynamics, enabling the model to learn diverse motion properties and better generalize across tasks.
@@
Within-Person Dynamics in Time-series: Variations in time-series signals can be related to both within-user dynamics and between-user dynamics. Comparing subsequences within a given user while learning the embedding can enable the learning of important within-user dynamics, which we ablate in w/o Sampling Within-Subject in Table 5. Here, we no longer draw candidates from sampling within the user, across time. Instead, our candidates are now only the other members within the batch, as well as a augmented version of the anchor, which mimics the candidates available to SimCLR. Afterwards, we follow the standard RelCon procedure, ranking the candidates and applying the RelCon loss. We see that this ablation results in a 4.49% drop in subsequence-level classification (AHMS-Subseq) by 4.49%, and by an even larger 7.25% for the workout-level. This demonstrates the importance of capturing within-user dynamics over time in order to achieve better performance for the longer-scale workout-level classifications. This intuition is reflected in Table 2, in which SimCLR, which does not capture within-user interactions, has relatively stronger performance in the subsequence-level metrics compared to the workout-level metrics.
Relative Rankings Soften the Impact of False Negatives/Positives: REBAR and RelCon each use a learnable distance to act as a static pseudo-labeling function to identify positives and negatives. However, out of all of the possible candidates, REBAR only identifies one candidate as positive and all others as negative. This unfortunately leads itself to false positives and false negatives. A false positive would occur when our distance measure identifies a candidate as positive, even when it originates from a different class than the anchor. False negatives occurs when there are multiple positive candidates within the pool of candidates, but only one of them is chosen to be a positive and the rest to be negative. This is an issue even across methods. SimCLR draws its negatives from the randomly chosen batch, and there is no guarantee that items in the batch are all semantically distinct from the anchor. Fortunately, RelCon’s learning approach allows us to capture how “positive” a candidate is relative to the other candidates, thereby lessening the effect of false pseudo-labels. In Tables 1 and 2, although both REBAR and RelCon used the same exact learned distance function, RelCon’s relative contrastive loss approach led to consistently stronger results, demonstrating the usefulness of relative positive rankings.
@@
Refining our Learned Distance: Table 5 shows that if we remove the reversible instance normalization (RevIN) component or replace the SparseMax with the traditional softmax operator within our learnable distance measure, then performance of our RelCon approach suffers. Removal of RevIN in particular causes a 11.66% drop in the velocity regression correlation, which shows the importance of capturing relative magnitude for predicting velocity.
Velocity DST   AHMS-Subseq AHMS-Workout    HHAR PAMAP2 MLP
↑ Corr ↑ Corr ↑ F1 ↑ F1 ↑ F1 ↑ F1
RelCon 0.8431 0.7559 38.56 55.28 57.63 85.38
w/o Augmentations -5.42% -13.88% -12.5% -17.93% -8.92% -15.12%
w/o RevIN -11.66% -6.01% -3.81% -4.11% -1.61% -4.65%
w/o SparseMax -0.87% -3.7% -5.06% -7.11% -4.34% -0.62%
w/o Sampling Within-Subject -1.4% -1.4% -4.49% -7.25% 2.24% -0.93%
Table 5: Ablations of RelCon components across tasks. The decrease in performance upon applying ablations shows the importance of key components of the RelCon approach. The large decrease in performance when augmentations are removed from the metric learning stage highlights the importance of learning a motif-similarity metric that is robust to invariances.
@@ 6 Conclusions
RelCon FM embeddings have strong performance across diverse tasks including motion velocity, double support time, and workout classification and across datasets, as compared to benchmarked approaches and models. This consistently strong performance of the frozen embeddings with simple probes shows that RelCon embeddings capture fundamental properties of motion relevant across tasks. To capture these motion properties, we introduced a number of novel modifications to our learned motif-based distance learning, including introducing augmentations to learn sensor invariances and using sparsemax to increase precision. We contribute a methodology for using the learned distance measure with a relative contrastive loss, in order to learn fine-grained interactions between and within users. Our RelCon foundation model showed the strongest, consistent performance in across all evaluated tasks and datasets, highlighting the value of the proposed approach.
@@ 7 Reproducibility
We will release the code for our RelCon foundation model publicly upon acceptance. This will include the RelCon training methodology, architecture, as well as the reproducible evaluation code from our “Benchmarking Evaluation” task, which utilizes public datasets. This can be used by the broader research community by providing code to easily re-train the RelCon approach in addition to providing a unified benchmarking task to guide model development.
@@ Appendix A Appendix
A.1 Model Implementation Details
@@
Aug Pred: We follow the approach described by Yuan et al. (2024), following the implementation located here: https://github.com/OxWearables/ssl-wearables?tab=readme-ov-file. Then we use a batch-size of 64 and train for 1e5 steps.
RelCon: We use the augmentations utilized by the aforementioned SimCLR model, batch-size of 64, temperature of 1, candidate set size of 20, and train for 1e5 steps.
@@
A.2 Extra AHMS Classification Results
Figure 5: Confusion Matrices for AHMS Classification at the Workout Level. We can see RelCon has the best performance. Unlike REBAR and AugPred, RelCon can better predict Outdoor running from Indoor Running. RelCon is also able to better predict Stair Climbing unlike the others.
@@ A.4 Justification for Using Single-Sensor-Accelerometer-Only Data
A.5 Elaborating on Table 1
Negatives in a SimCLR-based well-studied problem, with many recent works proposing novel methodologies to address this (Huynh et al., 2022; Jin et al., 2023; Chien & Chen, 2024). Additionally, the accelerometry SimCLR we are benchmarking (Tang et al., 2020) makes no distinction to model within-user interactions, by treating every subsequence as independent, and hence does not model within-user interactions. Both SimCLR and augmentation prediction creates instances from the original sequence via augmentations, and so they will be resistant to False Positives.
@@ A.7 Evaluation Dataset Descriptions
A.7.1 Comparison to a Large-Scale Pre-trained Accel Model (Yuan et al., 2024)
Yuan et al. (2024) has released their code publicly here, which contains exact data split and generations: https://github.com/OxWearables/ssl-wearables
Opportunity: 4-fold leave-one-subject-out cross validation with the sitting, standing, walking, and lying labels.
PAMAP2: 9-fold leave-one-subject-out cross validation with the lying, sitting, standing, walking, ascending stairs, descending stairs, vacuum cleaning, and ironing classes. This is the wrist-specific accelerometry data.
@@
Although Haresamudram et al. (2022) has not released their code publicly, we contacted the authors directly to ensure that we matched their exact splits and classes evaluated. Table 5.2 in our paper is then constructed by drawing from Table 3 and 4 in the original paper. Specifically, HHAR is drawn from Table 4, PAMAP2 from Table 3/4, and MotionSense from Table 3 (Haresamudram et al., 2022). Note that in Haresamudram et al. (2022), Augmentation Prediction is referred to as ”Multi-Task Self Supervision”.
HHAR: 5-fold leave-subject-out cross validation with the bike, sit, stairs down, stairs up, stand, and walk classes.
Motionsense: 5-fold leave-subject-out cross with the downstairs, upstairs, sitting, standing, walking, and jogging labels.
PAMAP2: 5-fold leave-subject-out cross validation with the lying, sitting, standing, walking, running, cycling, nordic walking, ascending stairs, descending stairs, vacuum cleaning, ironing, and rope jumping classes. This is the ankle-specific accelerometry data.
@@
A.9 Comparisons of RelCon vs. Softer or Harder Loss Functions
Velocity DST AHMS-Subseq AHMS-Workout
↑ Corr ↑ Corr ↑ F1 ↑ F1
RelCon 0.8431 0.7559 38.56 55.28
w/ Softer Metric Loss (Kim et al., 2019) -3.64% -6.11% -1.45% 1.39%
Table 6: Comparisons of RelCon vs. Softer or Harder Loss Functions
Our relative contrastive loss in Eq. 7 is particularly interesting because the relative contrastive loss is able to have a nice balance of softness. If we increase the softness in our loss function by utilizing a metric learning loss function (Kim et al., 2019), then this will hurt performance on the gait metric regression tasks. However, if we increase the hardness to a binary contrastive loss (i.e. REBAR), then this hurts performance across both types of tasks. Please see the table below, and we have added discussion to Section A.9.
A.10 Comparisons of RelCon vs. Yuan et al. (2024) with Same Encoder Backbone
We have re-trained our RelCon FM with the ResNet-18 backbone with a final encoding dimensionality of 1024, matching Yuan et al. (2024), and we show the results in the Table 7 below. Out of the 3/4 evaluations, RelCon continues to have stronger performance.
Opportunity (Wrist → Wrist) PAMAP2 (Wrist → Wrist)
Architecture Eval Method Pre-train Data ↑ F1 ↑ Kappa ↑ F1 ↑ Kappa
RelCon FM ResNet-18 Fine-tuned AHMS 98.1 ± 0.1 97.6 ± 1.2 97.9 ± 0.5 97.5 ± 0.8
Yuan et al. (2024)’s FM ResNet-18 Fine-tuned UKBioBank 59.5 ± 8.5 47.1 ± 10. 78.9 ± 5.4 76.9 ± 5.9
RelCon FM ResNet-18 MLP Probe AHMS 65.4 ± 12. 55.7 ± 12. 62.5 ± 13. 53.4 ± 6.6
Yuan et al. (2024)’s FM ResNet-18 MLP Probe UKBioBank 57.0 ± 7.8 43.5 ± 9.2 72.5 ± 5.4 71.7 ± 5.7
Table 7: Comparisons of RelCon vs. Yuan et al. (2024) with Same Encoder Backbone

### v5

@@ 5.3 Ablation Study Results
The Importance of Motion Augmentations: Invariances can capture conceptually important information for motion data, and augmentations have been used extensively in prior approaches to exploit this (Yurtman & Barshan, 2017; Florentino-Liaño et al., 2012; Zhong & Deng, 2014). For example, learning a 3D rotation invariance makes embeddings robust to how users wear their devices. RelCon uses these augmentations while learning the distance function, so that the distance function can identify instances that represent similar characteristics but differ from semantic-preserving transformations. Removing these augmentations from the distance function learning significantly degrades performance (>12% drop in 3/4 tasks, as shown in w/o Augmentations in Table 5), because the distance function fails to compare and identify semantically-similar instances.
@@
Within-Person Dynamics in Time-series: Variations in time-series signals stem from both within-user and between-user dynamics. Learning embeddings by comparing subsequences within the same user helps capture important temporal dynamics, as shown in the w/o Sampling Within-Subject ablation in Table 5. In this ablation, we no longer draw candidates from the within-user, across time. Instead the candidates are drawn from other batch members and augmented versions of them, similar to SimCLR. This leads to a 4% drop in subsequence-level classification (AHMS-Subseq) and a 7% drop in workout-level classification. These results highlight the importance of modeling within-user dynamics over time, especially for longer-scale workout-level tasks. This is also reflected in SimCLR’s performance in Table 2: SimCLR does not capture within-user interactions and also performs worse at the workout-level compare to the subsequence-level task
@@
Refining our Learned Distance: Table 5 shows that if we remove the reversible instance normalization, RevIN, component or replace the SparseMax with the traditional softmax operator within our distance function, then performance suffers. Removal of RevIN causes a 12% drop in the velocity correlation, which shows the importance of capturing relative magnitude for predicting velocity.
Relative Contrastive Loss is a Good Balance of “Softness”: RelCon’s relative contrastive loss is a softer version of REBAR’s hard binary contrastive loss. Both use a learnable distance to identify positives and negatives, but REBAR selects only one positive candidate, labeling the rest as negative, increasing the risk of false positives and negatives (Xu et al., 2024). False positives occur when a candidate is incorrectly identified as positive despite belonging to a different class, while false negatives arise when multiple valid positive candidates exist but only one is selected. The softer relative contrastive loss captures the relative “positiveness” of candidates, reducing the impact of false pseudo-labels. As shown in w/ Harder Binary Contrastive Loss in Table 5, RelCon’s relative contrastive loss improves performance >5% compared to the REBAR binary contrastive loss.
Alternatively, the relative loss function could be softened further with a metric loss function that directly learns the distances in the embedding space to be proportional to distances from the learned motif-based distance function. The distance function is now treated as an exact ground truth. In w/ Softer Metric Loss in Table 5, we replace the loss with the log-ratio metric loss, ${\ell(\textbf{X}_{\textrm{anc}})=\sum_{\textbf{X}_{\textrm{pos}},\textbf{X}_{\textrm{neg}}}\big{(}\log(\frac{\|E(\textbf{X}_{\textrm{anc}})-E(\textbf{X}_{\textrm{pos}})\|_{2}}{\|E(\textbf{X}_{\textrm{anc}})-E(\textbf{X}_{\textrm{neg}})\|_{2}})-\log(\frac{d(\textbf{X}_{\textrm{anc}},\textbf{X}_{\textrm{pos}})}{d(\textbf{X}_{\textrm{anc}},\textbf{X}_{\textrm{neg}})})\big{)}^{2}}$ from Kim et al. (2019). This results in a performance drop across all tasks, particularly in gait metric regression. This implies that a softer metric loss too exactly matches the distance function, obscuring the subtle features needed to differentiate gait. We aim for a balanced approach that allows the distance function to flexibly capture semantic groups, without strictly constraining the embedding space.
Stride Velocity Double Support Time Subseq-level Activity Workout-level Activity
↑ Corr ↑ Corr ↑ F1 ↑ F1
RelCon 0.8431 0.7559 38.56 55.28
w/o Augmentations -5.42% -13.88% -12.5% -17.93%
w/o RevIN -11.66% -6.01% -3.81% -4.11%
w/o SparseMax -0.87% -3.7% -5.06% -7.11%
w/o Sampling Within-Subject -1.40% -1.40% -4.49% -7.25%
w/ Softer Metric Loss (Kim et al., 2019) -3.64% -6.11% -1.45% 1.39%
Table 5: Ablations of Key RelCon Components. Removing augmentations in our distance learning approach results in a large drop in performance, highlighting the importance of learning a motif-based distance that is robust to accel-specific invariances. RevIN strongly affects the regression tasks. Within-subject interactions and SparseMax are important for classification. Finally, our relative loss strikes a good balance between hardness of the pos/neg psuedo-labels in the loss function.
@@ 6 Conclusions
We introduce RelCon, the first foundation model for accelerometry data with state-of-the-art performance on multiple diverse downstream motion tasks. RelCon embeddings exhibit consistently strong performance using only simple probes for diverse tasks including stride velocity, double support time, and human activity recognition, on multiple datasets. This demonstrates RelCon’s ability to capture fundamental properties of motion that are relevant across tasks. We achieve this through a novel motif-based distance learning approach that incorporates augmentations to learn sensor invariances, along with a relative contrastive loss approach for learning fine-grained interactions between- and within-subjects. Our results show that RelCon is highly effective for motif-based time-series data, such as accelerometry, and could potentially be effective for other quasi-periodic biosignals.
@@ 7 Reproducibility
The code for our RelCon foundation model can be found at https://github.com/maxxu05/relcon. This includes the RelCon training methodology, architecture, as well as the reproducible evaluation code from our “Benchmarking Evaluation” task, which utilizes public datasets. This can be used by the broader research community by providing code to easily re-train the RelCon approach in addition to providing a unified benchmarking task to guide model development.
8 Acknowledgments
Thank you to Di Wu, James Ochs, Sunny Chow, and Guillermo Sapiro at Apple for their support in this work. Special thanks as well to Harish Haresamudram for his help in replicating prior results, as well as Catherine Liu for her help in proofreading and editing.
Research reported here was supported by Apple and partially supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE-2039655 and the National Institutes of Health (NIH) under award P41EB028242. The opinions expressed in this article are the authors’ own and do not reflect the views of Apple, NSF, or NIH.
@@ Appendix A Appendix
A.1 Extra AHMS Field Human Activity Recognition Results
Figure 5: Confusion Matrices for AHMS Field Human Activity Recognition at the Workout Level. We can see RelCon has the best performance. Unlike REBAR and AugPred, RelCon can better predict Outdoor running from Indoor Running. RelCon is also able to better predict Stair Climbing unlike the others.
A.2 Model Implementation Details
@@
Augmentation Prediction: We follow the approach described by Yuan et al. (2024), following the implementation located here: https://github.com/OxWearables/ssl-wearables?tab=readme-ov-file. Then we use a batch-size of 64 and train for 1e5 steps.
@@
RelCon: We use the augmentations utilized by the aforementioned SimCLR model, batch-size of 64, temperature of 1, candidate set size of 20, and train for 1e5 steps. Please refer to our code at https://github.com/maxxu05/relcon for further details.
@@ A.4 Justification for Using Single-Sensor-Accelerometer-Only Data
A.5 Elaborating on the Table in Fig. 2
False Negatives in a SimCLR-based approach is a well-studied problem, with many recent works proposing novel methodologies to address this (Huynh et al., 2022; Jin et al., 2023; Chien & Chen, 2024). Additionally, the accelerometry SimCLR we are benchmarking (Tang et al., 2020) makes no distinction to model within-user interactions, by treating every subsequence as independent, and hence does not model within-user interactions. Both SimCLR and AugPred creates instances from the original sequence via augmentations, and so they will be resistant to False Positives.
@@ A.7 Evaluation Dataset Descriptions
A.7.1 Comparison to a Prior Large-Scale Pre-trained Accel Model (Yuan et al., 2024)
Yuan et al. (2024) has released their code publicly here, which contains exact data split and generations: https://github.com/OxWearables/ssl-wearables. Our codebase uses their exact splits22footnotemark: 2. Table 3 in our paper is then constructed by drawing from Table 2 in their original paper. Note that in Yuan et al. (2024), MLP Probe is referred to as ”Fine-tuned self-supervised after ConV layers”. The used dataset details are as follows:
  • Opportunity (Roggen et al., 2010): 4-fold leave-one-subject-out cross validation with the sitting, standing, walking, and lying labels.
  • PAMAP2 (Reiss, 2012): 9-fold leave-one-subject-out cross validation with the lying, sitting, standing, walking, ascending stairs, descending stairs, vacuum cleaning, and ironing classes. This is the wrist-specific accelerometry data.
@@
Although Haresamudram et al. (2022) has not released their code publicly, we contacted the authors directly to ensure that we matched their exact splits and classes evaluated. Our codebase uses their exact splits22footnotemark: 2. Table 5.2 in our paper is then constructed by drawing from Table 3 and 4 in their original paper. Specifically, HHAR is drawn from Table 4, PAMAP2 from Table 3/4, and MotionSense from Table 3 (Haresamudram et al., 2022). Note that in Haresamudram et al. (2022), Augmentation Prediction is referred to as ”Multi-Task Self Supervision”. The used dataset details are as follows:
  • HHAR (Blunck et al., 2015): 5-fold leave-subject-out cross validation with the bike, sit, stairs down, stairs up, stand, and walk classes.
  • Motionsense (Malekzadeh et al., 2018): 5-fold leave-subject-out cross validation with the downstairs, upstairs, sitting, standing, walking, and jogging labels.
  • PAMAP2 (Reiss, 2012): 5-fold leave-subject-out cross validation with the lying, sitting, standing, walking, running, cycling, nordic walking, ascending stairs, descending stairs, vacuum cleaning, ironing, and rope jumping classes. This is the ankle-specific accelerometry data.